OpenStudy (dls):

A box contains 10 red marbles,20 blue marbles and 30 green marbles.5 marbles are drawn from the box,what is the probability that 1))all are blue 2))atleast one is green

OpenStudy (dls):

@hartnn

OpenStudy (anonymous):

You first have to know is probability percents or fractions?

OpenStudy (anonymous):

^wrong

OpenStudy (kropot72):

The question implies that more than one marble is drawn. How many marbles are drawn? Are the marbles drawn without replacement?

OpenStudy (anonymous):

some typo here, probably a cut and past problem

OpenStudy (dls):

5,fixed

OpenStudy (kropot72):

Can we assume that 5 marbles are drawn without replacement?

OpenStudy (dls):

yes

OpenStudy (yrelhan4):

number of cases when all are blue 10c5 number of cases=50c5 .. divide.

OpenStudy (yrelhan4):

20c5 *

OpenStudy (kropot72):

\[P(5\ blue)=\frac{\left(\begin{matrix}20 \\ 5\end{matrix}\right)\left(\begin{matrix}40 \\ 0\end{matrix}\right)}{\left(\begin{matrix}60 \\ 5\end{matrix}\right)}\]

OpenStudy (yrelhan4):

part 2-- 1- (prob of no green marbles)

OpenStudy (dls):

got it! thanks!

OpenStudy (anonymous):

all are blue \[\frac{20}{60}\times \frac{19}{59}\times \frac{18}{58}\times \frac{17}{57}\times \frac{16}{56}\]

OpenStudy (anonymous):

at least one is green compute the probability that none are green, subtract from 1

OpenStudy (kropot72):

\[P(zero\ green)=\frac{5!55!}{60!}\] \[P(at\ least\ one\ green)=1-\frac{5!55!}{60!}=?\]